A 185 lb jogger is running at 8 mph. As he nears the end of his run, he decides to sprint, increasing his speed to 18 mph. How does his kinetic energy change? (Note: Express your answer as a multiple, e.g. his kinetic energy is 3 times as much, ½ as much, etc.)
Answer #1

let m = 185 lb
v1 = 8 mph
v2 = 18 mph


KE1 = (1/2)*m*v1^2

KE2 = (1/2)*m*v2^2

KE2/KE1 = (v2/v1)^2

KE2/KE1 = (18/8)^2

KE2/KE1 = 81/16

KE2 = 5.06*KE1 <<<<<<<<---------Answer

or KE2 = (81/16)*KE1